OpenStudy (anonymous):

Use substitution to find ordered pair thats not a solution for 6x+3y=6 2y+x=2 1.(0,2) 2. (1,0) 3.(-2,0) 4.(2,-2)

OpenStudy (anonymous):

try put all the values given in ans. if not satisfies in both eq. then that pair is not a solution.

OpenStudy (anonymous):

they all do

OpenStudy (anonymous):

no.

OpenStudy (cj49):

try the 3rd option

OpenStudy (anonymous):

3 works

OpenStudy (anonymous):

i mean thats the answeer

OpenStudy (anonymous):

huh

OpenStudy (cj49):

the second equation is Y + 2X=4

OpenStudy (anonymous):

check it clearly...2y+x=2

OpenStudy (cj49):

oops sorry i made a few mistakes...

OpenStudy (cj49):

even the 3rd option dosent satisfy the condition

OpenStudy (cj49):

6(-2)+3(0)=-12

OpenStudy (anonymous):

the question is which is not the solution.so only 3rd option doesn't support both the equations.so it will be the answer.

OpenStudy (abmon98):

The equation of a line or curve is a rule of determining whether or not the point with co-ordinates (x,y) lies on the line.

OpenStudy (mathmath333):

u can also draw the graph of both lines and determine ur answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

4